9. A genetic experiment with peas resulted in one sample of offspring that consisted of 408 green peas and 160 yellow peas.

a. Construct a 95​% confidence interval to estimate of the percentage of yellow peas.

b. It was expected that​ 25% of the offspring peas would be yellow. Given that the percentage of offspring yellow peas is not​25%, do the results contradict​ expectations?

a. a. Construct a 95​% confidence interval. Express the percentages in decimal form.

____<p<____

​(Round to three decimal places as​ needed.)

Given that the percentage of offspring yellow peas is not​ 25%, do the results contradict​ expectations?

​No, the confidence interval includes​ 0.25, so the true percentage could easily equal​ 25%

​Yes, the confidence interval does not include​ 0.25, so the true percentage could not equal​ 25%
Answer #1

Solution :

Given that,

a) n = 408 + 160 = 568

x = 160

Point estimate = sample proportion = \hat p = x / n = 160 / 568 = 0.282

1 - \hat p = 1 - 0.282 = 0.718

At 95% confidence level

\alpha = 1 - 95%

\alpha =1 - 0.95 =0.05

\alpha/2 = 0.025

Z\alpha/2 = Z0.025 = 1.960

Margin of error = E = Z\alpha / 2 * \sqrt ((\hat p * (1 - \hat p )) / n)

= 1.96 (\sqrt((0.282 * 0.718) / 568)

= 0.037

A 95% confidence interval for population proportion p is ,

\hat p - E < p < \hat p + E

0.282 - 0.037 < p < 0.282 + 0.037

(0.245 < p < 0.319)

b) ​Yes, the confidence interval does not include​ 0.25, so the true percentage could not equal​ 25%